OpenStudy (anonymous):

If h(x) = (f ◦ g)(x) and h(x) = ^3sqrt(7x + 1), find f(x) and g(x).

OpenStudy (freckles):

can you identify the function inside of the cube root function?

OpenStudy (anonymous):

\[h(x) = \sqrt[3]{7x+1}\]

OpenStudy (freckles):

\[\sqrt[3]{7x+1} \\ \sqrt[3]{\text{ asking for thing that is inside here }}\]

OpenStudy (freckles):

if you can identity that inside function you will have g(x)

OpenStudy (anonymous):

Whaaat??

OpenStudy (freckles):

I'm asking you to tell me what expression(function) is inside the cube root?

OpenStudy (anonymous):

I did

OpenStudy (freckles):

no you told me whole function

OpenStudy (freckles):

the question here is asking you to state inside and outside function

OpenStudy (freckles):

I'm asking for part of this question so far

OpenStudy (freckles):

what is the inside expression

OpenStudy (anonymous):

7x + 1

OpenStudy (freckles):

 perfect 7x+1 is g(x) g(x)=7x+1